The föllowing information applies to the questions displayed below Delph Company uses a job-order costing system and has two manufacturing departments-Molding and Fabrication. The company provided the following estimates at the beginning of the year: Molding Fabrication Machine-hours Fixed manufacturing overhead costs Variable manufacturing overhead cost per machine-hour 28,000 66,000 300,000 $1,040,000 38,000 s 740,000 $ 5.70 5.70 During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs- Job D-70 and Job C-200. It provided the following information related to those two jobs Job D-70: Direct materials cost Direct labor cost Machine-hours Molding Fabrication Total $. 374,000 $. 325,000 699,000 $. 220,000 $. 170,000 390,000 22,000 6,000 28,000 Job C-200: Direct materials cost Direct labor cost Machine-hours Molding Fabrication Total $. 280,000梟250,000 530,000 $. 110,000梟250,000 360,000 6,000 32,000 38,000 Delph had no underapplied or overapplied manufacturing overhead during the year. Exercise 2-15 Part 1 Required: 1. Assume Delph uses a plantwide predetermined overhead rate based on machine-hours. a. Compute the plantwide predetermined overhead rate b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200 c. If Delph establishes bid prices that are 140% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200? d. What is Delphs cost of goods sold for the year? Complete the question by entering your answers in the tabs given below Required 1A Required 1B Required 1CRequired 1D Assume Delph uses a plantwide predetermined overhead rate based on machine-hours. Compute the plantwide predetermined overhead rate. (Round your answer to 2 decimal places.) Predetermined overhead rate per MH
Answer #1

Solution:

a.) Prepare the statement as shown below to find total overhead under each of the process.

Particulars Modifications Fabrications Total
Variable manufacturing overheads rate $5.70 $5.70
No. of machine hours 28000 38000 66, 000
Variable manufacturing overheads $159,600 $216,000 $375, 600
Fixed manufacturing overheads $740,000 $300,000 $1,040,000
Total overhead $899,600 $516,000 $1,415,600

Overheads rate = total manufacturing/overheads number of machine hours

=1415600/66000 =$21.44

The predetermined manufacturing overheads rate is $21.44 per hour.

b.) Manufacturing overheads costs assigned to JOBD-70= Total manufacturing overheads×machine hours used in job total machine hours used

=1,415,600×28,000/66000=$600,557.57

Manufacturing overheads costs assigned to JOBC-200

=1,415,600×38,000/66000=$815,042.42

Total cost assignet to JOBD-70

=$699,000+$390,000+$600,557.57

=$1,689,557.57

Total cost assigned to JOBC-200

=$530, 000+$360, 000+$815, 042.42

=$1,705, 042.42

C.) The bid price for JOBD-70

=140% of $1,689, 557.57

=$2,365, 380.59

The bid price for JOBC-200

=140%of $1,705,042.42

=$2, 387,059.38

d. )

Cost of goods sold = beginning inventory + direct material + direct labor + overheads - closing inventory

=$0+($699, 000+$530, 000) +($390, 000+$360, 000) +$1, 415,600 -$0

=$3,394,600

The cost of goods sold is $3,394,600.